# Breaking better: Imperfections increase fracture resistance in architected lattices

### v1

@@ 4.2 Linking crack path morphology to work to failure
To systematically analyze this effect, we introduce a simple geometric measure of crack path tortuosity, namely, the maximum crack path deviation $\Delta_{y}$, defined as the difference between the maximum and minimum $y$-positions along the crack path (Fig. 5a). By examining the correlation between the work to failure and $\Delta_{y}$, we find that configuration A specimens exhibit a consistent and statistically significant increase in crack path deviation (Fig. 5b). This observation supports the interpretation that the deviations in crack trajectory induced by the imperfections contribute to increased work to failure. In contrast, configuration B shows only a slight, statistically insignificant increase in $\Delta_{y}$, suggesting that other mechanisms may play a more prominent role in the increased work to failure for this configuration (Fig. 5b); a point that will be discussed in the following section.
@@ 5 Discussion
A key observation in this study is the increased work to failure in imperfect lattice specimens and its potential link to deviations in the crack path. In configuration A, the presence of missing struts consistently led to a more tortuous crack path, as reflected by the increased in its maximum deviation. These deviations may influence the work to failure via two mechanisms. First, they increase the effective fracture surface relative to the crack length, thereby requiring more energy to propagate the crack. Second, the modified crack trajectory may locally reduce the stress intensity at the crack tip, delaying further propagation and requiring additional work. Both effects could contribute to the observed increase in work to failure. These findings indicate that crack path modifications, induced by local imperfections, serve as a mechanism to delay or extend failure in architected materials.
While configuration A exhibited a clear correlation between crack path deviation and increased work to failure, the behavior of configuration B suggested that other mechanisms may also contribute. In these specimens, the increase in work to failure was not accompanied by a statistically robust increase in crack path deviation. One possible explanation is the presence of crack bridging [42], where some struts within the crack path remain intact for part of the failure process, temporarily sustaining load and increasing energy dissipation (Fig. 5c). Another potential mechanism is crack pinning or temporary arrest at or near the location of the defect, which could delay propagation and result in a more extended loading phase [43, 44]. These mechanisms remain hypotheses, as they are not directly confirmed by the current analysis, but they are consistent with the observed mechanical response and warrant further investigation.
@@
In the reported experiments, the presence of a small number of missing struts did not affect the location or mechanism of crack initiation. In all tested specimens, both regular and imperfect, fracture consistently initiated at the notch root, where the stress concentration is highest. This observation indicates that, for the specific configurations considered herein, local imperfections placed away from the notch root do not significantly interfere with the initiation process. However, this result is not necessarily general. Some imperfections located closer to the notch root could modify the local stress field and potentially influence the onset of failure. Moreover, as the number of imperfections increases, the likelihood of one occurring near the notch becomes higher, which could compromise the robustness of crack initiation. These considerations highlight the need for further investigation into how the position and density of imperfections influence the early stages of fracture.
The results of this study suggest that local imperfections – typically seen as detrimental – can, in some cases, enhance fracture performance. In particular, the increased work to failure in imperfect specimens shows that small disruptions to the lattice topology can alter crack propagation while preserving strength and crack initiation. This observation points to the potential of defect engineering as a strategy to improve toughness in architected materials, echoing ideas explored in bioinspired and damage-tolerant designs where irregularities are used to guide or deflect cracks [45]. The present findings support the notion that crack path control, through deliberate confined geometric variation, may serve as a useful design principle for improving robustness in lattice-based structures.
This study focused on two specific imperfection configurations, both involving a small number of missing struts placed away from the notch root. While this choice allows for isolating and interpreting individual effects, the conclusions do not capture the full range of possible imperfection geometries or densities. In addition, the metric used to quantify crack path tortuosity is simple and does not account for the curvature or spatial complexity of the fracture trajectory. More refined geometric or energetic measures may reveal additional mechanisms or distinctions between configurations. Future work could explore a broader set of defect types, positions, and numbers to assess more systematically their influence on the fracture behavior. Combining such experimental studies with simulations could help disentangle local stress redistributions, crack pinning, and bridging effects. Extending the approach to different lattice topologies or loading modes would also clarify how general these observations are and inform design strategies for tougher, more failure-resistant architected materials.
@@ 6 Conclusion
In this work, we investigated the fracture behavior of architected materials with a triangular lattice topology by introducing controlled missing-strut imperfections and analyzing both macroscopic and local failure responses. Using a DIC-based approach, we tracked the crack path and identified the crack tip at the end of the test, providing detailed insight into local failure mechanisms. Our results showed that a small number of imperfections did not affect crack initiation or peak load but could significantly increase the work to failure. This increase was linked to deviations in the crack path, particularly in one configuration where the crack became more tortuous. In other cases, mechanisms such as crack bridging or pinning may have contributed to enhanced work to failure.
These findings suggest that local imperfections, often seen as defects, can be leveraged to improve energy dissipation during fracture. More broadly, they highlight crack path control as a potential design strategy in architected materials. Future studies should explore a wider range of imperfection types and placements, and integrate experiments with simulations to uncover the full potential of defect engineering for fracture-resistant design.
@@ 7 Acknowledgements
The authors acknowledge support from the Swiss National Science Foundation under the SNSF starting Grant (TMSGI2_211655). A.S. was supported by the Swiss National Science Foundation under grant number 200021_200343. The authors thank Dr. Mohit Pundir, Luca Michel, and Dr. Jan Van Dokkum for fruitful discussions and feedback.
@@ 10 Data Availability
Appendices
Appendix A Mesh backtracking
Figure 4c shows a triangular mesh, referred to as the DIC mesh, which precisely aligns with the strut joint locations of the lattice. This mesh was directly generated from the STL file used for 3D-printing. To position it, a backtracking procedure was applied, which was previously introduced in Ref. [46].
The process begins with the nominal STL geometry, as shown in Fig. 1. A fine triangular mesh was then constructed based on the location of the strut joints in the STL model, and cropped to match the selected region of interest for DIC analyses (Fig. 4c). From this mesh, a mask was generated in which pixels belonging to the mesh had their gray level equal to the mean gray level of the lattice and the other pixels had their gray levels set to the mean level of the image background. This mask corresponded to the nominal configuration in which the two meshes were created. Next, a coarser auxiliary mesh enclosing the region of interest was constructed. DIC was performed between the reference image of the experiment and the mask of the nominal configuration. Once convergence was reached, the auxiliary mesh was deformed according to the estimated displacement field and expressed in the nominal frame. Knowing the nodal displacements of the auxiliary mesh, the displacement of any other point (here the centers of the strut joints) from the reference to the nominal configurations was determined via mesh interpolation. The DIC mesh was finally backtracked by applying the opposite of these interpolated displacements to its nodes, thus mapping it from the nominal to the reference configuration.

### v2

@@ 5 Linking the crack path to the work to failure
By examining the correlation between the work to failure and the crack path tortuosity ($\Delta_{y}/h$), it is found that the higher crack path deviation observed for configuration A specimens consistently correlates with enhanced work to failure compared to regular triangular lattices (Fig. 6a). This observation supports the interpretation that the deviations in crack trajectory induced by the imperfections contribute to increased work to failure [37, 46]. Conversely, configuration B shows only a slight, statistically insignificant, increase in $\Delta_{y}/h$, thereby suggesting that other mechanisms may contribute to the increased work to failure for this configuration (Fig. 6b).
The work to failure gain for configuration B specimens correlates with the higher number of bridging struts during crack propagation compared to regular specimens (Fig. 6b). Conversely, the wide confidence interval characterizing configuration A specimens does not allow any effect of bridging to be distinguished for this configuration of imperfect lattices. These observations suggest that both crack path tortuosity and crack bridging act as toughening mechanisms in imperfect brittle lattices, with either mechanism prevailing depending on the specific location of the defect.
Figure 6: The work to failure ($W$) correlates with crack path tortuosity and bridging. $W$ as a function of (a) maximum tortuosity (normalized by the cell height) and (b) number of bridging struts, for regular and imperfect specimens. We normalize the work to failure with its elastic contribution ($W_{el}$). For all the measured quantities the error bars correspond to the 90% confidence interval.
@@ 6 Discussion
A key observation of this study is that the presence of a few vacancies alters the crack propagation path, thereby increasing the work to failure in imperfect lattice specimens. In configuration A specimens, the removal of struts consistently led to a more tortuous crack path, as reflected by the increase in its maximum deviation from the notch. These deviations may be the reason for higher work to failure as they increase the effective fracture surface relative to the crack length, thereby requiring more energy to propagate the crack [37, 47]. In configuration B specimens, the increase in work-to-failure correlates with the number of bridging struts. In analogy to the failure of composites [46], the simultaneous presence of load-bearing struts behind the crack tip is expected to contribute to greater crack growth resistance. These findings suggest that crack path modifications, induced by local imperfections, serve as a mechanism to delay failure in brittle architected materials.
Another potential toughening mechanism that we did not consider in this study is crack pinning or temporary arrest at or near the location of the defect, which could delay propagation and result in a more extended loading phase [48, 49, 50]. However, investigating this mechanism requires tracking the crack-tip advancement at each time step during testing, which is only possible by performing DIC analyses over a discrete lattice mesh and thus warrants further investigations.
@@
It is worth highlighting that, in our experiments, the presence of a small number of missing struts did not affect the location or mechanism of crack initiation. In all tested specimens, both regular and imperfect, fracture consistently initiated at the notch root, where the stress concentration is expected to be highest. This observation indicates that, for the specific six configurations considered herein, local imperfections placed one unit cell or more away from the notch root do not significantly interfere with the initiation process. This result is not necessarily general, as some imperfections located closer to the notch root could modify the local stress/strain fields and potentially influence the onset of failure.
A systematic sensitivity analysis examining how the increase in work to failure varies with the position of the removed strut, as well as with the number of removed struts, has not been performed owing to the complexity of the experimental procedures. A systematic variation of the placement of a single defect in spring network simulations [48] shows that a missing spring always increases or maintains the toughness (but never reduces it). However, the crack paths predicted by this spring network model do not agree with the paths observed experimentally, highlighting that more detailed simulation models are required to predict the failure of architected materials [31, 51, 52].
@@ 7 Conclusion
In this work, we investigated the impact of missing struts on the failure of pre-notched triangular lattices. Our results indicate that a few imperfections enhance the work to failure. Using a DIC-based approach to track crack propagation, we associated this improvement with increased path tortuosity and bridging. These findings experimentally demonstrate that introducing disorder to enhance toughness is also effective for brittle lattices. Future studies should explore a wider range of defect types and locations, and integrate experiments with simulations to fully exploit defect engineering for fracture-resistant design.
@@ 8 Acknowledgements
The authors acknowledge support from the Swiss National Science Foundation under the SNSF starting Grant (TMSGI2_211655). A.S. was supported by the Swiss National Science Foundation under grant number 200021_200343. The authors thank Dr. Mohit Pundir, Luca Michel, Flavio Lorez, Dr. Jan Van Dokkum, Dr. Leo De Waal, and Prof. Marcelo Dias for fruitful discussions and feedback.